OpenStudy (anonymous):

Graph the function. Compare the graph with the graph of f(x) = x Problem: h(x) = 6 +x Help?!

OpenStudy (anonymous):

h(x) = f(x) + 6

OpenStudy (anonymous):

Do you know about vertical shifts?

OpenStudy (anonymous):

so essentially h(x) = f(x) shifted 6 units above the origin

OpenStudy (anonymous):

The teacher been teaching us how to graph functions...no I do not.

OpenStudy (anonymous):

f(x) = x is a just a line through the origin

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh I see, so just one line goes through the origin, and then you plot the other point?

Directrix (directrix):

y = x goes through the origin because if x = 0, then y = 0. (0,0) are the coordinates of the origin. Two points determine a line. So, to draw the line y = x, you can use another point that satisfies the equation y = x. (1,1) is on the line y =x as is (2,2), (-3,-3) and so on. When you plot those, you get the graph shown above.

Directrix (directrix):

Plot the graph of y = x + 6 on the same set of axes as the first graph. Get two points on the line, say (0,6) and (-2, 4). Just choose an x and add 6 to get the y for that x. The lines have the same slope and will be parallel.

Directrix (directrix):

Courtesy of the "Wolf," the graphs are attached.

Directrix (directrix):

See the "Wolf" in person at the following link: http://tinyurl.com/akvmyb8